# Evolution of Concepts in Language Model Pre-Training

### v1

@@ Abstract
Language models obtain extensive capabilities through pre-training. However, the pre-training process remains a black box. In this work, we track linear interpretable feature evolution across pre-training snapshots using a sparse dictionary learning method called crosscoders. We find that most features begin to form around a specific point, while more complex patterns emerge in later training stages. Feature attribution analyses reveal causal connections between feature evolution and downstream performance. Our feature-level observations are highly consistent with previous findings on Transformer’s two-stage learning process, which we term a statistical learning phase and a feature learning phase. Our work opens up the possibility to track fine-grained representation progress during language model learning dynamics.
@@ 1 Introduction
Pre-training (Radford et al., 2018; Devlin et al., 2019) has emerged as the dominant paradigm for developing frontier large language models (LLM) (OpenAI, 2023; Grattafiori et al., 2024; Yang et al., 2025). Despite its remarkable success, the pre-training process remains largely a black box. While scaling laws (Hestness et al., 2017; Kaplan et al., 2020; Bahri et al., 2021) reveal the predictable relationships between compute, data, and loss, they offer limited insight into the internal reorganization occurring within the model parameters. Other theoretical frameworks on learning dynamics, including neural tangent kernel (Jacot et al., 2018), information bottleneck theory (Tishby & Zaslavsky, 2015; Shwartz-Ziv & Tishby, 2017), and singular learning theory (Watanabe, 2009; Lau et al., 2024; Wang et al., 2025a) — provide high-level explanations on why we can observe generalization and grokking (Power et al., 2022; Nanda et al., 2023). However, a fundamental question — how do LLMs actually develop their capabilities internally during pre-training — has remained largely opaque.
Recent advances in mechanistic interpretability, particularly through dictionary learning methods based on sparse autoencoders (SAEs), have begun to illuminate the internal representation of neural networks (Bricken et al., 2023; Templeton et al., 2024; Gao et al., 2025). By disentangling the phenomenon of superposition (Elhage et al., 2022), sparse autoencoders show their capabilities in extracting millions of features from LLM activations, demonstrating that LLMs encode human-interpretable concepts as linear directions in their activation spaces. However, these analyses have predominantly focused on studying fully-trained models. Consequently, the process of how features initially emerge and evolve throughout the training process remains largely unexplored.
In this paper, we propose to track feature evolution across pre-training snapshots using crosscoders (Lindsey et al., 2024), a variant of sparse autoencoders designed to simultaneously identify and align features from a family of correlated model activations. While originally introduced to resolve cross-layer superposition and track features distributed across layers, we adapt this approach to analyze activations from different training checkpoints. By applying cross-snapshot crosscoders, we can observe where features emerge, rotate, and degenerate, thereby providing deeper insight into the internal dynamics of model pre-training. Our main contributions are as follows:
@@ 2 Related Works
Multiple theoretical frameworks explain neural network training dynamics. Neural Tangent Kernel (NTK) theory (Jacot et al., 2018) establishes that infinite-width networks evolve as kernel machines with fixed kernels during gradient descent, with extensions to finite-width corrections (Dyer & Gur-Ari, 2020) and modern architectures (Yang, 2020; Yang & Littwin, 2021). Recent work identifies phase transitions and lazy regimes in training dynamics (Kumar et al., 2024; Zhou et al., 2025). Information Bottleneck (IB) theory (Tishby & Zaslavsky, 2015) formulates deep learning as optimizing compression-prediction tradeoffs. Shwartz-Ziv & Tishby (2017) empirically demonstrates two distinct training phases—initial fitting followed by compression—which aligns with our findings in Section 6. Singular Learning Theory (SLT) treats neural networks as singular statistical models. Watanabe (1999; 2009) introduces the Real Log Canonical Threshold to provide geometric complexity measures predicting phase transitions, with recent advances enabling practical estimation at scale through Local Learning Coefficients (Lau et al., 2024; Furman & Lau, 2024; Wang et al., 2025a). Our work complements these theoretical frameworks by providing detailed mechanistic accounts of feature evolution during transformer pre-training, bridging high-level theory with empirical observations.
@@
The superposition hypothesis posits that models use linear representations (Bengio et al., 2014; Alain & Bengio, 2017; Vargas & Cotterell, 2020) to embed more features than neurons (Olah et al., 2020; Elhage et al., 2022). Sparse Autoencoders (SAEs) (Bricken et al., 2023; Huben et al., 2024; He et al., 2024) extract monosemantic features from superposition using sparse dictionary learning, with subsequent improvements and scaling efforts (Gao et al., 2025; Rajamanoharan et al., 2024). Ge et al. (2024) and Dunefsky et al. (2024) propose transcoders, an SAE variant that predicts future activations to improve circuit tracability. Recently, Lindsey et al. (2024) introduces crosscoders to simultaneously read and write to multiple layers. Mishra-Sharma et al. (2024); Minder et al. (2025) leverages crosscoders for capturing the difference between pretrained models and their chat-tuned versions.
@@ 3 Tracking the Evolution of Features
A cross-snapshot crosscoder (Lindsey et al., 2024) operates on the parameterized function family $\mathcal{A}$ over corpus $\mathcal{C}$ (Figure 1). The crosscoder architecture is defined by:
@@
We use Pythia-160M and Pythia-6.9B (Biderman et al., 2023) for our experiments throughout this work. Pythia is a Transformer language model suite with well-controlled training settings and accessibility to training snapshots. We select the middle layers (Layer 6 of Pythia-160M and Layer 16 of Pythia-6.9B) for training crosscoders. To balance training cost with the granularity of feature evolution analysis, we select 32 snapshots out of 154 open-source snapshots ranging from step 0 to 143,000, with a stratified sampling approach: (1) all 20 snapshots before step 10,000 to capture early feature evolution with maximum temporal resolution, and (2) 12 evenly spaced snapshots from later training stages, containing 4 snapshots from steps 14,000–34,000 and 8 snapshots from steps 47,000–143,000. We use SlimPajama (Shen et al., 2023), a comprehensive text dataset covering a variety of data sources to sample activations.
@@ Results.
Figure 2 demonstrates crosscoder performance in decomposing model activations into sparse feature representations. We evaluate reconstruction quality (measured by activation variance explained) and sparsity (measured by L0 norm averaged across snapshots). Increasing dictionary size $n_{\text{features}}$ yields significant Pareto improvements across both metrics. Our crosscoders demonstrate comparable results to per-snapshot SAEs (Appendix A).
@@ 4.1 Overview of Feature Evolution
    Initialization features that exist from random initialization, exhibit a sudden drop and recovery around step $128$, then gradually decay. The existence of these features has been established by Bricken et al. (2023) and Heap et al. (2025).
@@
Previous studies suggest that loss curves comprise discrete phase changes, each contributing to different circuits at distinct training stages, with evidence from toy model features (Elhage et al., 2022), 5-digit addition (Nanda et al., 2023), and in-context learning (Olsson et al., 2022). From a feature evolution perspective, we investigate whether features emerge abruptly enough to constitute the fundamental units of these phase transitions. By evaluating the time spent by each feature from initial emergence to peakness, we find that emergence steepness exhibits considerable variation, demonstrating the coexistence of both gradually-formed and abruptly-appearing features.
@@ 4.2 Correlation between Emergence Step and Complexity
We follow Cunningham & Conerly (2024) to score feature complexity for 100 randomly sampled emergent features based on their top activations, ranging from 1 (simple) to 5 (complex). The result in Figure 5(a) reveals a non-trivial correlation with moderate strength between peak timing and complexity scores, suggesting that more complex features tend to emerge later in training. The scoring rubrics and complete prompt for automated complexity scoring can be found in Appendix B.
@@ 4.3 Case Study on Typical Cross-Snapshot Features
We employ simple rule-based approaches to find several well-studied feature types in a crosscoder with 32,768 features trained on Pythia-6.9B, including previous token features (which activate based on the preceding token), induction features (which activate on the second [A] in patterns [A][B]…[A][B] and help predict [B]), and context-sensitive features. These feature categories are extensively documented in prior research on neurons and SAE features (Gurnee et al., 2024; Huben et al., 2024; Ge et al., 2024).
Figure 5(b) demonstrates the distinct temporal pattern in feature emergence: previous token features arise early (around steps 1,000–5,000), while induction and context-sensitive features appear later and over a wider range (mostly steps 10,000–100,000). This suggests a general emergence hierarchy, from previous token to induction to context-sensitive features, which aligns with their increasing complexity. This finding is also consistent with the causal dependency between induction heads and previous token heads (Olsson et al., 2022).
@@ 5 Connecting Microscopic Evolution to Macroscopic Behaviors
One of the ambitious missions of mechanistic interpretability research is to connect feature-level findings with the model’s downstream performance. To this end, we employ attribution-based circuit tracing techniques (Syed et al., 2023; Ge et al., 2024; Marks et al., 2025) to investigate the causal effects of crosscoder features formation on downstream task metrics improvements.
@@
Let metric $m:\mathbb{R}^{d_{\text{model}}}\to\mathbb{R}$ be an arbitrary scalar-valued function of model activations $a^{\theta}(x)$. [^1] The original notation in Marks et al. (2025) uses $m$ as a function of input $x$, where the computation graph flows through nodes of interest. Since we focus on feature effects within a single layer (e.g., layer 6 of Pythia-160M), we simplify $m$ as a function of model activations to avoid complex intervention notation. To quantify the causal effect of each feature activation $f_{i}(x)$ on metric $m$, we first decompose model activations into per-feature representations:
@@
For structured downstream tasks with clean/corrupted input pairs, such as subject-verb agreement (Finlayson et al., 2021), we can employ the full framework of attribution patching (Syed et al., 2023; Marks et al., 2025) by emphasizing differences between clean and corrupted inputs:
@@
We evaluate our method on subject-verb agreement (SVA) (Finlayson et al., 2021), induction, and indirect object identification (IOI) (Wang et al., 2023) tasks, using 1000 samples each to identify critical features.
@@ 6 Observations of A Statistical-to-Superposition Transition
This two-stage structure is deeply consistent with the fitting-to-compressing phase transition predicted by the information bottleneck theory (Shwartz-Ziv & Tishby, 2017).
@@
Following previous work on language model learning statistical patterns (Takahashi & Tanaka-Ishii, 2017; Xu et al., 2019; Choshen et al., 2022; Belrose et al., 2024), we compute the KL divergence between the true token distribution $Q$ and the model’s predicted distribution $P$. We randomly sample 10M tokens from SlimPajama to approximate both distributions. We then evaluate unigram KL divergence $\mathrm{D}_{\mathsf{KL}}(P(x)\parallel Q(x))$ and bigram KL divergence $\mathrm{D}_{\mathsf{KL}}(P(x_{i}\mid x_{i-1})\parallel Q(x_{i}\mid x_{i-1}))$.
The results show that both unigram and bigram KL divergences rapidly converge to low values during early training (Figure 7(a) and 7(b)). Furthermore, the training losses during this period approach the theoretical minimum achievable if the model perfectly matched the true token distributions—i.e., the entropy of these distributions. This suggests that the model primarily learns to fit statistical regularities (Zipf’s law (Zipf, 1935; Piantadosi, 2014)) during the early training stage, which explains the dense nature of internal representations at this phase.
@@
To directly measure superposition status at each snapshot, we adapt the approach from Elhage et al. (2022) and compute the dimensionality of each feature at snapshot $\theta$ as:
@@
where $\hat{W}_{\text{dec},j}^{\theta}$ is the normalized version of decoder vector $W_{\text{dec},j}^{\theta}$. In contrast to its original application in toy models with ground-truth features, we apply this metric to crosscoder features due to their consistent alignment across training snapshots, providing insight into superposition dynamics.
@@ 7 Limitations
Superposition has been proven to be a general phenomenon in deep neural networks (Elhage et al., 2022). However, our analysis focuses on feature evolution in Pythia suite models using their open-source training snapshots. While previous research on feature universality (Wang et al., 2025b) suggest our method might be able to generalize to different architectures, datasets, post training dynamics, and tasks beyond language modeling, the extent to which feature evolution patterns are consistent across diverse settings remains to be established. We leave broader generalization studies to future work.

### v2

@@ Abstract
Language models obtain extensive capabilities through pre-training. However, the pre-training dynamics remains a black box. In this work, we track linear interpretable feature evolution across pre-training snapshots using a sparse dictionary learning method called crosscoders. We find that most features begin to form around a specific point, while more complex patterns emerge in later training stages. Feature attribution analyses reveal causal connections between feature evolution and downstream performance. Our feature-level observations are highly consistent with previous findings on Transformer’s two-stage learning process, which we term a statistical learning phase and a feature learning phase. Our work opens up the possibility to track fine-grained representation progress during language model learning dynamics. Our code is available at https://github.com/OpenMOSS/Language-Model-SAEs.
@@ 1 Introduction
Pre-training (radford2018gpt; devlin2019bert) has emerged as the dominant paradigm for developing frontier large language models (LLM) (openai2024gpt4; grattafiori2024llama3; yang2025qwen3). Despite its remarkable success, the pre-training process remains largely a black box. While scaling laws (hestness2017scaling; kaplan2020scalinglaws; bahri2021explainscalinglaws) reveal the predictable relationships between compute, data, and loss, they offer limited insight into the internal reorganization occurring within the model parameters. Other theoretical frameworks on learning dynamics, including neural tangent kernel (jacot2018ntk), information bottleneck theory (tishby2015informationbottleneck; shwartz2017information), and singular learning theory (watanabe2009slt; lau2024llc; wang2025rllc) — provide high-level explanations on why we can observe generalization and grokking (power2022grokking; nanda2023grokking). However, a fundamental question — how do LLMs actually develop their capabilities internally during pre-training — has remained largely opaque.
Recent advances in mechanistic interpretability, particularly through dictionary learning methods based on sparse autoencoders (SAEs), have begun to illuminate the internal representation of neural networks (bricken2023monosemanticity; templeton2024scalingmonosemanticity; gao2025scalingsae). By disentangling the phenomenon of superposition (elhage2022superposition), sparse autoencoders show their capabilities in extracting millions of features from LLM activations, demonstrating that LLMs encode human-interpretable concepts as linear directions in their activation spaces. However, these analyses have predominantly focused on studying fully-trained models. Consequently, the process of how features initially emerge and evolve throughout the training process remains largely unexplored.
In this paper, we propose to track feature evolution across pre-training snapshots using crosscoders (lindsey2024crosscoder), a variant of sparse autoencoders designed to simultaneously identify and align features from a family of correlated model activations. While originally introduced to resolve cross-layer superposition and track features distributed across layers, we adapt this approach to analyze activations from different training checkpoints. By applying cross-snapshot crosscoders, we can observe where features emerge, rotate, and degenerate, thereby providing deeper insight into the internal dynamics of model pre-training. Our main contributions are as follows:
@@ 2 Related Works
Multiple theoretical frameworks explain neural network training dynamics. Neural Tangent Kernel (NTK) theory (jacot2018ntk) establishes that infinite-width networks evolve as kernel machines with fixed kernels during gradient descent, with extensions to finite-width corrections (dyer2020asymptotics) and modern architectures (yang2020tensorprograms3; yang2021tensorprograms2b). Recent work identifies phase transitions and lazy regimes in training dynamics (kumar2024grokkingntk; zhou2025twophase). Information Bottleneck (IB) theory (tishby2015informationbottleneck) formulates deep learning as optimizing compression-prediction tradeoffs. shwartz2017information empirically demonstrates two distinct training phases—initial fitting followed by compression—which aligns with our findings in Section 6. Singular Learning Theory (SLT) treats neural networks as singular statistical models. watanabe1999slt; watanabe2009slt introduces the Real Log Canonical Threshold to provide geometric complexity measures predicting phase transitions, with recent advances enabling practical estimation at scale through Local Learning Coefficients (lau2024llc; furman2024llc; wang2025rllc). Our work complements these theoretical frameworks by providing detailed mechanistic accounts of feature evolution during transformer pre-training, bridging high-level theory with empirical observations.
@@
The superposition hypothesis posits that models use linear representations (bengio2014representationlearning; alain2017linearprobe; vargas2020linear) to embed more features than neurons (olah2020zoom; elhage2022superposition). Sparse Autoencoders (SAEs) (bricken2023monosemanticity; hubun2024sae; he2024llamascope) extract monosemantic features from superposition using sparse dictionary learning, with subsequent improvements and scaling efforts (gao2025scalingsae; rajamanoharan2024jumprelu). ge2024automatically and dunefsky2024transcoders propose transcoders, an SAE variant that predicts future activations to improve circuit tracability. Recently, lindsey2024crosscoder introduces crosscoders to simultaneously read and write to multiple layers. mishra2024crosscoder; minder2025crosscoders leverages crosscoders for capturing the difference between pretrained models and their chat-tuned versions.
@@ 3 Tracking the Evolution of Features
A cross-snapshot crosscoder (lindsey2024crosscoder) operates on the parameterized function family $\mathcal{A}$ over corpus $\mathcal{C}$ (Figure 1). The crosscoder architecture is defined by:
@@
We use Pythia-160M and Pythia-6.9B (biderman2023pythia) for our experiments throughout this work. Pythia is a Transformer language model suite with well-controlled training settings and accessibility to training snapshots. We select the middle layers (Layer 6 of Pythia-160M and Layer 16 of Pythia-6.9B) for training crosscoders. To balance training cost with the granularity of feature evolution analysis, we select 32 snapshots out of 154 open-source snapshots ranging from step 0 to 143,000, with a stratified sampling approach: (1) all 20 snapshots before step 10,000 to capture early feature evolution with maximum temporal resolution, and (2) 12 evenly spaced snapshots from later training stages, containing 4 snapshots from steps 14,000–34,000 and 8 snapshots from steps 47,000–143,000. We use SlimPajama (shen2023slimpajama), a comprehensive text dataset covering a variety of data sources to sample activations.
@@ Results.
Figure 2 demonstrates crosscoder performance in decomposing model activations into sparse feature representations. We evaluate reconstruction quality (measured by activation variance explained) and sparsity (measured by L0 norm averaged across snapshots). Increasing dictionary size $n_{\text{features}}$ yields significant Pareto improvements across both metrics.
To examine whether crosscoders can extract cross-snapshot features and align identical features in consistent directions within the feature space, we compare crosscoder performance against corresponding SAEs trained individually on each Pythia snapshot. We train SAEs using identical settings and hyperparameters as crosscoders on each snapshot. Figures 9(a) and 9(b) demonstrate that crosscoders achieve comparable performance in L0 norm and explained variance at each snapshot, even when SAEs are optimized for individual snapshots where they should have a theoretical advantage.
@@
Potential failure modes of crosscoder feature alignment.
Before assessing crosscoder features, we further discuss two theoretical concerns about cross-snapshot feature alignment:
Will crosscoders misalign unrelated features? Such misalignment would strongly contradict the optimization objective of crosscoders. The distinct activation patterns of unrelated features would conflict when forced into the same dimension. Once a feature activates, it would cause others to activate simultaneously, introducing noise in the reconstruction. In terms of results, misalignment would lead to the polysemanticity of crosscoder features, damaging the consistency of their interpretation, which our interpretability evaluation demonstrates does not occur.
Will crosscoders split shared features? Suppose features from different snapshots represent the same underlying concept with identical activation patterns, splitting them across multiple dimensions would be suboptimal: it wastes representational capacity in feature space, as both previous works (gao2025scalingsae; templeton2024scalingmonosemanticity) and Appendix A.4 prove that feature space dimensionality is crucial for better reconstruction fidelity. Nevertheless, in rare cases, we do observe feature splitting among highly active features. These features share semantic and directional similarity but exhibit subtly different activation patterns that justify separate dimensions. We detail this phenomenon in Appendix F.
@@ 4.1 Overview of Feature Evolution
    Initialization features that exist from random initialization, exhibit a sudden drop and recovery around step $128$, then gradually decay. The existence of these features has been established by bricken2023monosemanticity and heap2025saerandom.
@@
Previous studies suggest that loss curves comprise discrete phase changes, each contributing to different circuits at distinct training stages, with evidence from toy model features (elhage2022superposition), 5-digit addition (nanda2023grokking), and in-context learning (olsson2022induction). From a feature evolution perspective, we investigate whether features emerge abruptly enough to constitute the fundamental units of these phase transitions. By evaluating the time spent by each feature from initial emergence to peakness, we find that emergence steepness exhibits considerable variation, demonstrating the coexistence of both gradually-formed and abruptly-appearing features.
@@ 4.2 Correlation between Emergence Step and Complexity
We follow cunningham2024complexity to score feature complexity for 100 randomly sampled emergent features based on their top activations, ranging from 1 (simple) to 5 (complex). The result in Figure 6(a) reveals a non-trivial correlation with moderate strength between peak timing and complexity scores, suggesting that more complex features tend to emerge later in training. The scoring rubrics and complete prompt for automated complexity scoring can be found in Appendix B.
@@ 4.3 Case Study on Typical Cross-Snapshot Features
We employ simple rule-based approaches to find several well-studied feature types in a crosscoder with 32,768 features trained on Pythia-6.9B, including previous token features (which activate based on the preceding token), induction features (which activate on the second [A] in patterns [A][B]…[A][B] and help predict [B]), and context-sensitive features. These feature categories are extensively documented in prior research on neurons and SAE features (gurnee2024universalneuron; hubun2024sae; ge2024automatically).
Figure 6(b) demonstrates the distinct temporal pattern in feature emergence: previous token features arise early (around steps 1,000–5,000), while induction and context-sensitive features appear later and over a wider range (mostly steps 10,000–100,000). This suggests a general emergence hierarchy, from previous token to induction to context-sensitive features, which aligns with their increasing complexity. This finding is also consistent with the causal dependency between induction heads and previous token heads (olsson2022induction).
@@ 5 Connecting Microscopic Evolution to Macroscopic Behaviors
One of the ambitious missions of mechanistic interpretability research is to connect feature-level findings with the model’s downstream performance. To this end, we employ attribution-based circuit tracing techniques (syed2023attributionpatching; ge2024automatically; marks2025sparsefeaturecircuits) to investigate the causal effects of crosscoder features formation on downstream task metrics improvements.
@@
Let metric $m:\mathbb{R}^{d_{\text{model}}}\to\mathbb{R}$ be an arbitrary scalar-valued function of model activations $a^{\theta}(x)$. [^1] The original notation in marks2025sparsefeaturecircuits uses $m$ as a function of input $x$, where the computation graph flows through nodes of interest. Since we focus on feature effects within a single layer (e.g., layer 6 of Pythia-160M), we simplify $m$ as a function of model activations to avoid complex intervention notation. To quantify the causal effect of each feature activation $f_{i}(x)$ on metric $m$, we first decompose model activations into per-feature representations:
@@
For structured downstream tasks with clean/corrupted input pairs, such as subject-verb agreement (matthew2021sva), we can employ the full framework of attribution patching (syed2023attributionpatching; marks2025sparsefeaturecircuits) by emphasizing differences between clean and corrupted inputs:
@@
We evaluate our method on subject-verb agreement (SVA) (matthew2021sva), induction, and indirect object identification (IOI) (wang2023ioi) tasks, using 1000 samples each to identify critical features.
@@ 6 Observations of A Statistical-to-Superposition Transition
This two-stage structure is deeply consistent with the fitting-to-compressing phase transition predicted by the information bottleneck theory (shwartz2017information).
@@
Following previous work on language model learning statistical patterns (takahashi2017statistical; xu2019frequency; choshen2022grammar; belrose2024statistics; svete2024ngram; nguyen2024ngram; chen2024jetexpansionsngram), we compute the KL divergence between the true token distribution $Q$ and the model’s predicted distribution $P$. We randomly sample 10M tokens from SlimPajama to approximate both distributions. We then evaluate unigram KL divergence $\mathrm{D}_{\mathsf{KL}}(P(x)\parallel Q(x))$ and bigram KL divergence $\mathrm{D}_{\mathsf{KL}}(P(x_{i}\mid x_{i-1})\parallel Q(x_{i}\mid x_{i-1}))$.
The results show that both unigram and bigram KL divergences rapidly converge to low values during early training (Figure 8(a) and 8(b)). Furthermore, the training losses during this period approach the theoretical minimum achievable if the model perfectly matched the true token distributions—i.e., the entropy of these distributions. This suggests that the model primarily learns to fit statistical regularities (Zipf’s law (zipf1935psychobiology; piantadosi2014zipfs)) during the early training stage, which explains the dense nature of internal representations at this phase.
@@
To directly measure superposition status at each snapshot, we adapt the approach from elhage2022superposition and compute the dimensionality of each feature at snapshot $\theta$ as:
@@
where $\hat{W}_{\text{dec},i}^{\theta}$ is the normalized version of decoder vector $W_{\text{dec},i}^{\theta}$. In contrast to its original application in toy models with ground-truth features, we apply this metric to crosscoder features due to their consistent alignment across training snapshots, providing insight into superposition dynamics.
@@ 7 Limitations
Superposition has been proven to be a general phenomenon in deep neural networks (elhage2022superposition). However, our analysis focuses on feature evolution in Pythia suite models using their open-source training snapshots. While previous research on feature universality (wang2025universality) suggest our method might be able to generalize to different architectures, datasets, post training dynamics, and tasks beyond language modeling, the extent to which feature evolution patterns are consistent across diverse settings remains to be established. We leave broader generalization studies to future work.